# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ Abstract
Recent advancements in machine learning have improved performance while also increasing computational demands. While federated and distributed setups address these issues, their structure is vulnerable to malicious influences. In this paper, we address a specific threat, Byzantine attacks, where compromised clients inject adversarial updates to derail global convergence. We combine the trust scores concept with trial function methodology to dynamically filter outliers. Our methods address the critical limitations of previous approaches, allowing functionality even when Byzantine nodes are in the majority. Moreover, our algorithms adapt to widely used scaled methods like Adam and RMSProp, as well as practical scenarios, including local training and partial participation. We validate the robustness of our methods by conducting extensive experiments on both synthetic and real ECG data collected from medical institutions. Furthermore, we provide a broad theoretical analysis of our algorithms and their extensions to aforementioned practical setups. The convergence guarantees of our methods are comparable to those of classical algorithms developed without Byzantine interference.
@@ 1  Introduction
As the field of machine learning expands, researchers are confronted with challenges stemming from more complex models, larger computational demands, and data privacy. To address these issues, distributed and federated learning scenarios are among the most popular that have been developed (Kairouz et al., 2021; Konečný et al., 2016; Li et al., 2020). These approaches are crucial for a wide range of tasks (Smith et al., 2017; McMahan et al., 2017; Verbraeken et al., 2020), yet they introduce several complications. Making learning process multi-node though leads to various threats, especially those related to data storage and transmission. These vulnerabilities can manifest as device malfunctions, incorrect data relays, or even initial data corruption (Biggio et al., 2012). Moreover, (Wang et al., 2023) demonstrated how sophisticated adversarial attacks on data integrity pose significant challenges, fundamentally compromising the effectiveness of the learning process. In the midst of them are Byzantine attacks. They occur in networks where certain workers, known as Byzantines, may deliberately or accidentally introduce corrupted data, disrupting the learning process (Blanchard et al., 2017; Yin et al., 2018; Karimireddy et al., 2021b). This underscores the critical need for robust defenses.
@@
This paper specifically examines the threat of Byzantine attacks and highlights the limitations of existing protection mechanisms, particularly their reliance on strong assumptions. In response, we propose an approach that leverages diverse concepts to develop a universal method, free from these constraints and applicable to practical scenarios.
Related work. Methods resistant to Byzantine attacks are currently crucial for solving optimization problems. Classical existing methods for distributed optimization, such as SGD (Robbins & Monro, 1951; Bottou, 2012; Recht & Ré, 2013; McMahan et al., 2017), Adam (Kingma & Ba, 2014; Reddi et al., 2020) and Scaffold (Karimireddy et al., 2020), share a basic aggregation rule: averaging the gradients or models received from devices. However, they cease to operate as soon as even a single Byzantine worker appears in the network (Blanchard et al., 2017).
@@
Given the critical importance of this problem, numerous publications addressed it (Feng et al., 2014; Damaskinos et al., 2019). Initial approaches proposed robust aggregation rules for data from devices, such as Coordinate-wise median, Trimmed Mean (Yin et al., 2018), Krum (Blanchard et al., 2017), Bulyan (Mhamdi et al., 2018). However, sophisticated attacks, such as ALIE (Baruch et al., 2019) or Inner Product Manipulation (Xie et al., 2020), managed to circumvent these aggregation rules by shifting the mean they seek to find.
@@
Moreover, these rules are non-robust even in the absence of attacks, for example, in the case of imbalanced classes. This issue was addressed in (Karimireddy et al., 2021b), where Centered Clip (CC) technique was revealed. Additionally, the authors highlighted that the aforementioned Byzantine-robust methods cannot converge with any predetermined accuracy. Given the importance of this issue, they added client momentum, effectively tackling this problem. Another approach to combat the Byzantines is the application of the variance reduction technique, originally developed to eliminate irreducible errors in stochastic methods. It was proposed as an effective mean of mitigating the presence of noise in computing stochastic gradient estimates (Gorbunov et al., 2023). Later this idea was developed in (Malinovsky et al., 2023). These methods demonstrated significant improvements in resilience against Byzantine attacks. However, they also contain some notable limitations. First, variance reduction methods exhibit moderate convergence in deep learning applications and are prone to overfitting (Defazio & Bottou, 2019). Additionally, all aforementioned approaches suffer from a serious shortcoming: they require the majority of devices to be honest.
@@
Another approach to achieving solutions with any specified accuracy involves techniques like validation tests (Alistarh et al., 2018; Allen-Zhu et al., 2020) or computation checks (Gorbunov et al., 2022). Nevertheless, they still require a majority of honest devices and rely on strict assumptions in analysis (Alistarh et al., 2018; Allen-Zhu et al., 2020).
@@
While much work on the Byzantines focused on the distributed case with homogeneous data, a different series of papers proposed methods for more realistic training scenarios by allowing data heterogeneity (Wu et al., 2020; El-Mhamdi et al., 2021; Data & Diggavi, 2021; Nguyen et al., 2022). This corresponds, for example, to the federated setting. Methods were primary built around robust aggregation (Karimireddy et al., 2021a; Chang et al., 2019; Data & Diggavi, 2021; Allouah et al., 2024c; Dorfman et al., 2024; Allouah et al., 2024b), and variance reduction techniques (Allouah et al., 2023). One of the most advanced approaches assigned coefficients to clients based on their reliability, using these trust scores to perform gradient steps (Cao et al., 2020; Yan et al., 2024). These studies provided a foundation for Byzantine-robust optimization in the federated setup, but they still suffered from previously mentioned drawbacks.
To circumvent the requirement for a majority of honest workers, several methods attempted to use ground truth data on the server to filter out compromised updates from workers. Thus, Zeno algorithm (Xie et al., 2019), required a validation dataset on the server to evaluate each incoming gradient by comparing it with this dataset. We refer to the function computed on such a validation dataset as the trial function. Nonetheless, Zeno uses this concept only partially. It calculates trust scores using trial functions but then combines the results mostly by regular averaging, except for devices with low trust scores. This results in a significant dependency on the choice of trusted devices. Later, (Cao & Lai, 2019) proposed an alternative algorithm that filters compromised updates. It does this by comparing the updates with a noisy gradient approximation computed on a small dataset, which is effectively equivalent to a validation dataset in Zeno. However, classical averaging techniques, which discard clients with low trust scores, are limited by their high sensitivity to hyperparameters. Specifically, they depend on threshold values for determining malicious devices.
@@
The authors of (Xie et al., 2019; Cao & Lai, 2019) extended their results to handle data heterogeneity by requiring the server to have a representative sample of all device data. However, this assumption is unrealistic in real-world scenarios, undermining the fundamental achievements of federated learning regarding privacy. Moreover, similar approaches (Guo et al., 2021, 2024) also accumulate user data on the server, casting doubt on their applicability.
@@
In addition to these limitations of methods in heterogeneous scenarios, many studies in this field are predominantly heuristic and lack rigorous theoretical analysis (Yan et al., 2024; Guo et al., 2021, 2024; Chang et al., 2019; Xu & Lyu, 2020; Rodríguez-Barroso et al., 2022; Nguyen et al., 2022; Zhang et al., 2022; Huang et al., 2024). In addition, in some studies, the practical component seems to be flawed due to the absence of experiments assessing test accuracy (Gorbunov et al., 2023). Furthermore, theoretical frameworks often do not align with practical part, such as in (Cao et al., 2020), where homogeneous data sampling is assumed despite focusing on the federated learning.
@@
Furthermore, when addressing problems in the distributed or federated setups, local methods (Woodworth et al., 2020; Khaled et al., 2020; Gorbunov et al., 2021b; Nguyen et al., 2022), as well as the partial participation scenario (Yang et al., 2021; Kairouz et al., 2021; Sadiev et al., 2022; Nguyen et al., 2022), is typically assumed. While these options improve computational efficiency and reduce data transmission overhead, only a few studies (Data & Diggavi, 2021; Malinovsky et al., 2023; Allouah et al., 2024a; Dorfman et al., 2024) address these aspects, whereas the majority of works do not. In addition, research is often limited to SGD-like methods, neglecting adaptive algorithms such as Adam (Kingma & Ba, 2014) and RMSProp (Tieleman & Hinton, 2012), which are widely used in machine learning. Such gaps can lead to significant discrepancies between theoretical expectations and practical implementations, impacting the reliability of the results.
Given the challenges and gaps identified in the existing literature, we aim to advance trust scores methodology and trial function concept to enhance the applicability of defense mechanisms.
Contributions. Our main results in tackling Byzantines can be summarized as follows.
$\bullet$ Combine trust scores with the trial function approach. The trial loss function is based on a subset of the training sample stored on the server. To derive trust scores, we evaluate how the gradients sent from devices minimize the trial loss. Two methods embody the ideas of our approach:
  • $(a)$
@@
    The first method assigns weights for the gradients sent from each device based on the extent to which these gradients reduce the trial function in each iteration. In real networks, honest stochastic gradients may increase the target loss. We account for this by incorporating weights from the previous epoch and a momentum parameter for a more stable convergence.
  • $(b)$
    The second approach employs a similar concept of selecting weights. At each step, an additional optimization problem is solved to determine optimal weights that minimize the trial function.
@@
$\bullet$ Milder assumptions. Unlike most existing studies, our approach requires only one reliably honest worker instead of a majority. Moreover, unlike previous trial function-based methods that assume data homogeneity (Cao et al., 2020; Gorbunov et al., 2022), our algorithms work under the more realistic assumption of data similarity in federated learning. In Byzantine literature, it is a common premise for ensuring convergence (Karimireddy et al., 2021a; Gorbunov et al., 2023; Yan et al., 2024).
@@
$\bullet$ Extensions. We adapt our algorithms to practically important scenarios often overlooked in research.
@@
    Partial participation. Our algorithms incorporate the option for partial participation. Thus, devices may not participate in every learning step, and the group involved in the attack can vary across iterations.
@@
  • $(d)$
    Finding scores from validation. We also introduce a method that employs the concept of a trial function in a slightly different manner. This approach focuses on the trial data stored on the server, where we conduct pairwise validation between the server and the device. Based on the degree of alignment between the predictions made by the device and those generated by the server, we assign weights to each device accordingly.
$\bullet$ Convergence guarantees. We prove upper bounds on convergence rates of main methods and extensions presented for the smooth problem under various assumptions regarding the convexity of the target function (strong convexity, convexity, non-convexity).
$\bullet$ Experiments. We conduct comprehensive experiments, including tests on real data, demonstrating the superiority of our method in both previously studied attacks and scenarios where other methods fail. Our experiments are performed on the CIFAR-10 dataset and real ECG data, using ResNet-18 and ResNet-1D18 neural networks, respectively.
2  Setup
To establish the groundwork for our study, we begin by defining the problem with the assumptions on which our work is based. We consider the problem that is often encountered in distributed machine learning with $\mathcal{D}_{i}$ being an unknown distribution of the training sample data on the $i$-th device:
We consider a setup involving $n$ workers connected to a central server. These workers are divided into two categories: good (or honest) workers, indicated by $\mathcal{G}$ and Byzantines, indicated by $\mathcal{B}$. At each iteration $t$, the sets $\mathcal{G}(t)$ and $\mathcal{B}(t)$ are redefined. This allows the composition of honest and Byzantine workers to vary dynamically over time. At every step, we assume that the set of honest workers $\mathcal{G}(t)$ is nonempty, i.e., $G(t)\overset{def}{=}|\mathcal{G}(t)|\geqslant 1$. The presence of at least one honest worker per iteration maintains the feasibility of the optimization. During training, we do not know the number of Byzantines at each iteration and use it only in theoretical analysis of the worst-case scenario.
To tackle Byzantine attacks, we introduce a pivotal component of our methodology — a trial loss function $\hat{f}$. In the homogeneous setting (1) with $\mathcal{D}_{i}=\mathcal{D}$, we take a separate sample from $\mathcal{D}$ but in a smaller volume than the entire data. The trial function calculated on this data forms $\hat{f}$. Under the heterogeneous data scenario, we sample from the distribution $\mathcal{D}_{1}$ on the server to obtain a delayed data for $\hat{f}$ (indexing the server does not violate generality; we further consider it a device with index 1). Formally, we can write the trial function as $\hat{f}(x)=\frac{1}{N}\sum_{i=1}^{N}f_{1}(x,\xi_{i}),$ where $N$ is the number of samples in $\hat{f}$. This function is stored on the server (obviously an honest device). The sample distribution of the trial function is similar to the entire distribution $\mathcal{D}$ due to the data similarity property. Besides, in practical scenarios, a server may not be able to share the entire dataset, providing only a sample of size $N$. Depending on the size of this sample, $f_{1}$ may differ from $\hat{f}$. Nevertheless, the larger the volume of this delayed sample is, the closer $\hat{f}$ approximates the function $f_{1}$ (discussed in Lemma C.1). Here we outline the assumptions under which we establish the convergence rates.
@@ 2  Setup
    $\mu$-strongly convex if it satisfies the inequality for all $x,y\in\mathbb{R}^{d}$:
    $\hat{f}(y)\geqslant\hat{f}(x)+\left\langle\nabla\hat{f}(x),y-x\right\rangle+\frac{\mu}{2}\|y-x\|^{2}.$
@@
    convex if it satisfies the inequality for all $x,y\in\mathbb{R}^{d}$:
    $\hat{f}(y)\geqslant\hat{f}(x)+\left\langle\nabla\hat{f}(x),y-x\right\rangle.$
@@
    non-convex if it has at least one (not necessarily unique) minimum, i.e.,
    $\hat{f}(\hat{x}^{*})=\inf\limits_{x\in\mathbb{R}^{d}}\hat{f}(x)>-\infty.$
@@
$\mathbb{E}\|g_{i}(x,\xi_{i})-\nabla f_{i}(x)\|^{2}\leqslant\sigma^{2}\quad\text{for all }x\in\mathbb{R}^{d}.$
@@
In over-parameterized models, introducing a positive $\delta_{2}$ can sometimes reduce the value of $\delta_{1}$. Several studies on Byzantine-resilient optimization have explored heterogeneous scenarios where honest workers handle distinct local functions (Wu et al., 2020; Karimireddy et al., 2021a). Thus, when heterogeneity is limited to $\delta_{2}$-bounded settings ($\delta_{1}=0$), achieving a predefined accuracy becomes infeasible in the presence of Byzantine workers, even with a trusted Parameter Server (Karimireddy et al., 2021a). In that way, we consider a more general assumption on heterogeneity.

### v5

# Bant: Byzantine Antidote via Trial Function and Trust Scores

@@ Abstract
Recent advancements in machine learning have improved performance while also increasing computational demands. While federated and distributed setups address these issues, their structures remain vulnerable to malicious influences. In this paper, we address a specific threat: Byzantine attacks, wherein compromised clients inject adversarial updates to derail global convergence. We combine the concept of trust scores with trial function methodology to dynamically filter outliers. Our methods address the critical limitations of previous approaches, allowing operation even when Byzantine nodes are in the majority. Moreover, our algorithms adapt to widely used scaled methods such as Adam and RMSProp, as well as practical scenarios, including local training and partial participation. We validate the robustness of our methods by conducting extensive experiments on both public datasets and private ECG data collected from medical institutions. Furthermore, we provide a broad theoretical analysis of our algorithms and their extensions to the aforementioned practical setups. The convergence guaranties of our methods are comparable to those of classical algorithms developed without Byzantine interference.
Codehttps://github.com/brain-lab-research/Bant
@@ 1 Introduction
As the field of machine learning expands, researchers are confronted with challenges stemming from increasingly complex models and larger computational demands. To address these issues, distributed and federated learning scenarios have been developed (kairouz2021advances; konevcny2016federated; li2020federated). These approaches are crucial for a wide range of tasks (smith2017federated; mcmahan2017communication; verbraeken2020survey), yet, they introduce several complications. Making the training process multi-node leads to threats related to data storage and transmission. These vulnerabilities manifest as device malfunctions, incorrect data relays, or even initial data corruption (biggio2012poisoning). (wang2023adversarial) demonstrated how adversarial attacks on data integrity compromise training. In the midst of them are Byzantine attacks. They occur in networks where certain workers, known as Byzantines, may corrupt data (blanchard2017machine).
@@
This paper specifically examines the threat of Byzantine attacks. We highlight the limitations of existing protection mechanisms, particularly their reliance on strong assumptions. In response, we propose an approach that leverages diverse concepts to develop a universal method, free from these constraints and applicable to practical scenarios.
Setup. We examine the problem often encountered in distributed machine learning, with $\mathcal{D}_{i}$ representing an unknown distribution of the training sample data on the $i$-th device:
We consider a setup involving $n$ workers connected to a central server. These workers are divided into two categories: Good (or Honest) workers, indicated by $\mathcal{G}$, and Byzantines, indicated by $\mathcal{B}$. At each iteration $t$, the sets $\mathcal{G}(t)$ and $\mathcal{B}(t)$ are redefined. This allows the composition of honest and Byzantine workers to vary dynamically over time. At every step, we assume that the set of honest workers $\mathcal{G}(t)$ is nonempty, i.e., $G(t)\mathrel{\mathop{:}}=|\mathcal{G}(t)|\geqslant 1$. During training, the number of Byzantines at each iteration remains unknown. This number is used only in the theoretical analysis of the worst-case scenario.
2 Related work
@@
Sophisticated attacks. Methods resistant to Byzantine attacks are crucial for solving optimization problems. Classical methods for distributed optimization, such as SGD (robbins1951stochastic; bottou2012stochastic; recht2013parallel; mcmahan2017communication), Adam (kingma2014adam; reddi2020adaptive), and Scaffold (karimireddy2020scaffold), average the received gradients or models. However, they cease to operate when even a single Byzantine worker appears in the network (blanchard2017machine). Given the critical importance of this problem, numerous publications have addressed it (feng2014distributed; damaskinos2019aggregathor). Initial approaches proposed robust aggregation rules for data from devices, such as Coordinate-wise median, Trimmed Mean (yin2018byzantine), Krum (blanchard2017machine), Bulyan (mhamdi2018hidden). However, sophisticated attacks, such as ALIE (baruch2019little) or Inner Product Manipulation (xie2020fall), managed to circumvent these aggregation rules by shifting the mean they sought to find.
@@
Clipping, Momentum, and Variance Reduction. Aggregation rules are non-robust even in the absence of attacks. For example, this occurs in cases of imbalanced classes. This issue was addressed in (karimireddy2021learning), where Centered Clip (CC) technique was revealed. Additionally, the authors highlighted that the aforementioned Byzantine-robust methods cannot converge with any predetermined accuracy. Given the importance of this issue, they added client momentum, effectively addressing the problem. Another approach to combat the Byzantines is the application of the variance reduction technique. Originally introduced to eliminate irreducible errors in stochastic methods, it was subsequently proposed as an effective way to mitigate the presence of noise in computing stochastic gradient estimators (gorbunov2023variance). Then this idea was developed in (malinovsky2023byzantine).
These methods demonstrated significant improvements in resilience against Byzantine attacks. However, they also contain notable limitations. First, variance reduction methods exhibit moderate convergence in deep learning applications and are prone to overfitting (defazio2019ineffectiveness). Additionally, all aforementioned approaches suffer from the requirement that the majority of devices must be honest.
@@
Validation tests. Another approach to achieving solutions with any specified accuracy involves techniques such as validation tests (alistarh2018byzantine; allen2020byzantine) or computation checks (gorbunov2022secure). Nevertheless, they still require a majority of honest devices and rely on strict assumptions in the analysis (alistarh2018byzantine; allen2020byzantine).
@@
Heterogeneous setup. While much work on the Byzantines focused on the distributed case with homogeneous data, a different series of papers allowed for data heterogeneity (wu2020federated; el2021collaborative; data2021byzantine; nguyen2022flame). This corresponds, for example, to the federated setting. Methods were primarily built around robust aggregation (karimireddy2021byzantine; chang2019cronus; data2021byzantine; allouah2024robust; dorfmandynamic; allouah2024boosting), and variance reduction techniques (allouah2023fixing). One of the most advanced approaches assigned coefficients (trust scores) to clients based on their reliability, using these scores to perform gradient steps (cao2020fltrust; yan2024recess). These studies provided a foundation for Byzantine-robust optimization in the federated setup, yet suffered from the aforementioned drawbacks.
Majority of attackers – trial function. To avoid requiring a majority of honest workers, several methods leverage server-held ground-truth data to filter compromised updates. Zeno (xie2019zeno) implements this idea by maintaining a small validation set on the server and scoring each incoming gradient against the gradient computed on that set. We define the function evaluated on the validation set as the trial function. As for Zeno, it uses the trial function to compute trust scores but primarily aggregates updates through simple averaging, down-weighting or excluding devices with very low trust. As a result, its performance depends strongly on the devices that are treated as trusted. (cao2019distributed) proposed an alternative that filters updates by comparing them to a noisy gradient approximation computed on a small dataset, which is effectively equivalent to using a validation set as in Zeno. The idea of using a small server-side dataset to validate local updates was also utilized in SageFlow (park2021sageflow), which handles majority adversarial clients. However, this method performs poorly under an attack ratio of 60%. Moreover, all aggregation schemes that discard clients based on trust scores remain sensitive to hyperparameters, as they require threshold choices to label devices as malicious. It requires tuning, and we discuss such instability regarding Zeno in Appendix, Table 7.
@@
The authors of (xie2019zeno; cao2019distributed; park2021sageflow) extended their results to address data heterogeneity by requiring the server to possess a representative sample of all device data. However, this assumption is unrealistic in real-world scenarios, undermining the fundamental achievements of federated learning with respect to privacy. Approaches (guo2021siren; guo2024siren+) also accumulate user data on the server, raising doubts about applicability.
@@
Other shortcomings of previous research. In addition to aforementioned limitations, many studies in this field are predominantly heuristic and lack rigorous theoretical analysis (yan2024recess; guo2021siren; guo2024siren+; chang2019cronus; xu2020reputation; rodriguez2022dynamic; nguyen2022flame; zhang2022fldetector; huangself). Moreover, in some studies, the practical component seems flawed due to the absence of experiments assessing test accuracy (gorbunov2023variance). Furthermore, theoretical settings often do not align with practical aspects. For instance, in (cao2020fltrust), homogeneous data sampling is assumed while focusing on the federated learning. Besides, the analysis of Sageflow is confined to an unrealistic strongly convex setup.
@@
Furthermore, when addressing problems in the distributed or federated setups, local methods (woodworth2020minibatch; khaled2020tighter; gorbunov2021local; nguyen2022flame), as well as the partial participation scenario (yang2021achieving; kairouz2021advances; sadiev2022decentralized; nguyen2022flame), is typically assumed. While these options improve computational efficiency and reduce data transmission overhead, only a few studies (data2021byzantine; malinovsky2023byzantine; allouahbyzantine; dorfmandynamic) address these aspects, whereas the majority of works do not. In addition, research is often limited to SGD-like methods, neglecting adaptive algorithms such as Adam (kingma2014adam) and RMSProp (tieleman2012lecture), which are widely used in machine learning. Given the challenges and gaps identified in the existing literature, we aim to advance the trust score methodology and trial function concept to enhance defense mechanisms.
Contributions
Our main results are summarized as follows.
@@
$\bullet$ Combine trust scores with the trial function approach. Trial loss is based on a subset of the training data stored on the server. Weights are assigned to the gradients sent from each device based on the extent to which these gradients reduce the trial function in each iteration. In real networks, honest stochastic gradients may increase the target loss. We account for this by incorporating weights from the previous epoch and a momentum parameter for more stable convergence.
@@
$\bullet$ Milder assumptions. Unlike most existing studies, our approach requires only one reliably honest worker instead of a majority. Moreover, unlike previous trial function-based methods that assume data homogeneity (cao2020fltrust; gorbunov2022secure), our algorithms operate under the more realistic assumption of data similarity in federated learning.
@@
$\bullet$ Extensions. We adapt our algorithms to important scenarios that are often overlooked in research.
@@
    Partial participation. Our algorithms incorporate the option for partial participation. Thus, devices may not participate in every learning step, and the attackers may vary across iterations.
@@
$\bullet$ Convergence guarantees. We prove upper bounds on the convergence rates of the main methods and extensions presented for the smooth problem under various assumptions regarding the convexity of the target function (strong convexity, convexity, non-convexity).
$\bullet$ Experiments. We demonstrate the superiority of our method in both previously studied attacks and scenarios where alternative methods fail. Our experiments are performed on the CIFAR-10 dataset and real ECG data, utilizing ResNet-18 and ResNet-1D18 neural networks, respectively. Additionally, we validate our approach to Learning-to-Rank tasks by training a Transformer-based ranking model.
3 Methodology
To tackle Byzantine attacks, we introduce a pivotal component of our methodology – a trial loss function $\hat{f}$. In the homogeneous setting (1) with $\mathcal{D}_{i}=\mathcal{D}$, we take a separate sample from $\mathcal{D}$, but in a smaller volume than the entire dataset. The trial function calculated on this data forms $\hat{f}$. Under the heterogeneous data scenario, we sample from the distribution $\mathcal{D}_{1}$ on the server to obtain delayed data for $\hat{f}$ (indexing the server does not violate generality; we further consider it a device with index 1). Formally, we can write the trial function as $\hat{f}(x)=\frac{1}{N}\sum_{i=1}^{N}f_{1}(x,\xi_{i}),$ where $N$ is the number of samples in $\hat{f}$. This function is stored on the server (obviously, an honest device). The sample distribution of the trial function is similar to the entire distribution $\mathcal{D}$ due to the property of data similarity. Besides, in practical scenarios, a server may not be able to share the entire dataset, providing only a sample of size $N$. Depending on the size of this sample, $f_{1}$ may differ from $\hat{f}$. Nevertheless, the larger the volume of this delayed sample, the closer $\hat{f}$ approximates the function $f_{1}$ (discussed in Lemma C.1 in Appendix). A small public or synthetic trial dataset is a practical assumption in Byzantine-robust federated learning. Methods such as FLTrust (cao2020fltrust) and Zeno (xie2019zeno) utilize such datasets. Here we outline the assumptions under which we establish the convergence rates.
@@ 3 Methodology
    $\mu$-strongly convex if for all $x,y\in\mathbb{R}^{d}$, it satisfies: ${\hat{f}(y)\geqslant\hat{f}(x)+\langle\nabla\hat{f}(x),y-x\rangle+\frac{\mu}{2}\|y-x\|^{2}.}$
@@
    convex if for all $x,y\in\mathbb{R}^{d}$, it satisfies: ${\hat{f}(y)\geqslant\hat{f}(x)+\langle\nabla\hat{f}(x),y-x\rangle.}$
@@
    non-convex if it has at least one (not necessarily unique) minimum, i.e., ${\hat{f}(\hat{x}^{*})=\inf\limits_{x\in\mathbb{R}^{d}}\hat{f}(x)>-\infty.}$
@@
$\mathbb{E}\|g_{i}(x,\xi_{i})-\nabla f_{i}(x)\|^{2}\leqslant\sigma^{2},\quad\text{for all }x\in\mathbb{R}^{d}.$
We also assume data similarity – a common premise for ensuring convergence in Byzantine literature (karimireddy2021byzantine; gorbunov2023variance; yan2024recess).
@@
In over-parameterized models, introducing a positive $\delta_{2}$ can sometimes reduce the value of $\delta_{1}$. Several studies have explored heterogeneous scenarios in which honest workers handle distinct local functions (wu2020federated; karimireddy2021byzantine). Note that achieving a predefined accuracy becomes feasible in the presence of Byzantines only when heterogeneity is limited to $\delta_{2}$-bounded settings $(\delta_{1}=0)$(karimireddy2021byzantine). Under a more general assumption, which we utilize, the term with $\delta_{1}$ inadvertently appears in the estimate.